OpenStudy (anonymous):

Find a number t such that the line passing through the two points has slope -2. (5, t); (-3, 4)

OpenStudy (anonymous):

\[\frac{ 4 - t }{ -3-5 }=-2\]\[\frac{ 4-t }{ -8 }=-2\]\[4-t=16\]\[t=-12\]

OpenStudy (anonymous):

Thank you!